# Differentially Private Bilevel Optimization: Efficient Algorithms with Near-Optimal Rates

### v1

@@ 5 Conclusion and discussion
We provided novel algorithms and lower bounds for differentially private bilevel optimization, with near-optimal rates for the convex setting and state-of-the-art rates for the nonconvex setting. There are some interesting open problems for future work to explore: (1a) What are the optimal rates for DP nonconvex bilevel ERM and SO? Since the optimal rates for standard single-level DP nonconvex ERM and SO are still unknown, a first step would be to answer: (1b) Can we match the SOTA rate for single-level non-convex ERM lowy2024make in BLO? Incorporating variance-reduction in DP BLO seems challenging. (2) This work was focused on fundamental theoretical questions about DP BLO, but another important direction is to provide practical implementations and experimental evaluations.
@@ Appendix A More privacy preliminaries
[…]
@@ B.1 Conceptual algorithms and excess risk upper bounds
Lemma B.2 (Utility Guarantee, [15, Corollary 1]).
@@
[…]
@@
[…]
@@ B.2 Generalization error of the regularized exponential mechanism for bilevel SCO
Another advantage of Regularized Exponential Mechanism is that it can have a good generalization error.
Lemma B.7 ([22]).
@@
$\mathbb{E}_{x\sim\pi_{Z},Z\sim P^{n}}[\Phi(x)-\Phi^{*}]\leq\frac{G^{2}}{\mu n}+\frac{d_{x}}{k}.$
Theorem B.8 (Re-statement of Theorem 3.3).
@@
Grant Assumption 2.1 and parts 3 and 4 of Assumption 2.2. Assume $\widehat{\Phi}_{Z}$ and $\Phi$ are convex for all $Z$. Sampling $\widehat{x}$ from a distribution proportional to $\exp(-k(\widehat{\Phi}_{Z}(x)+\mu\|x\|^{2}/2))$ with $k=O\left(\frac{\mu n^{2}\epsilon^{2}}{G^{2}\log(1/\delta)}\right)$ and $G=(L_{f,x}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+\frac{L_{g,y}\beta_{f,xy}}{\mu_{g}})$ is $(\epsilon,\delta)$-DP. Moreover,
@@
  • setting $\mu=\frac{G\sqrt{d_{x}\log(1/\delta)}}{nD_{x}\epsilon}$, we achieve excess risk
@@
  • setting $\mu=\frac{G\sqrt{d_{x}\log(1/\delta)}}{nD_{x}\epsilon}+\frac{G}{D_{x}\sqrt{n}}$, the population loss has the following guarantee:
    $\mathbb{E}\Phi(\widehat{x})-\Phi^{*}\leq O\left(\left(L_{f,x}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+\frac{L_{g,y}\beta_{f,xy}}{\mu_{g}}\right)D_{x}\left(\frac{\sqrt{d_{x}\log(1/\delta)}}{n}+\frac{1}{\sqrt{n}}\right)\right)$
@@
As for the population loss, with the setting of $\mu$ and $k$, by Lemma B.7, we have
$\mathbb{E}\Phi(\widehat{x})-\Phi^{*}\lesssim\frac{d_{x}}{k}+\frac{G^{2}}{\mu n}+\mu D_{x}^{2}=\frac{d_{x}G^{2}\log(1/\delta)}{\mu n^{2}\varepsilon^{2}}+\frac{G^{2}}{\mu n}+\frac{\mu D_{x}^{2}}{2}=O\left(GD_{x}(\frac{\sqrt{d_{x}\log(1/\delta)}}{n}+\frac{1}{\sqrt{n}})\right).$

### v2

@@ 5 Conclusion and discussion
We provided novel algorithms and lower bounds for differentially private bilevel optimization, with near-optimal rates for the convex setting and state-of-the-art rates for the nonconvex setting. There are some interesting open problems for future work to explore: (1) What are the optimal rates for DP bilevel convex SO? As discussed in Remark 3.5, we believe that it should be possible, though challenging, to obtain an improved $O(1/\sqrt{n})$ generalization error bound nearly matching the lower bound in Remark 3.11. (2a) What are the optimal rates for DP nonconvex bilevel ERM and SO? Since the optimal rates for standard single-level DP nonconvex ERM and SO are still unknown, a first step would be to answer: (2b) Can we match the SOTA rate for single-level non-convex ERM lowy2024make in BLO? Incorporating variance-reduction in DP BLO seems challenging. (3) This work was focused on fundamental theoretical questions about DP BLO, but another important direction is to provide practical implementations and experimental evaluations.
@@ Appendix A More privacy preliminaries
Let $\varepsilon\geq\ 0,\delta,\delta^{\prime}\in[0,1)$. Assume $\mathcal{A}_{1},\cdots,\mathcal{A}_{T}$, with $\mathcal{A}_{t}:\mathcal{Z}^{n}\times\mathcal{X}\to\mathcal{X}$, are each $(\varepsilon,\delta)$-DP  $\forall t=1,\cdots,T$. Then, the adaptive composition $\mathcal{A}(Z):=\mathcal{A}_{T}(Z,\mathcal{A}_{T-1}(Z,\mathcal{A}_{T-2}(X,\cdots)))$ is $(\varepsilon^{\prime},T\delta+\delta^{\prime})$-DP for
@@ B.1 Conceptual algorithms and excess risk upper bounds
Lemma B.2 (Utility Guarantee, [17, Corollary 1]).
@@
$=\int_{0}^{1}\nabla_{x}f(tx+(1-t)x^{\prime},\widehat{y}^{*}_{Z}(tx+(1-t)x^{\prime}),z_{i})dt\cdot(x-x^{\prime})$
$-\int_{0}^{1}\nabla_{x}f(tx+(1-t)x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(tx+(1-t)x^{\prime}),z_{i})dt\cdot(x-x^{\prime})$
@@
$\left\|\int_{0}^{1}[\nabla_{x}f(tx+(1-t)x^{\prime},\widehat{y}^{*}_{Z}(tx+(1-t)x^{\prime}),z_{i})-\nabla_{x}f(tx+(1-t)x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(tx+(1-t)x^{\prime}),z_{i})]dt\right\|$
$\leq\int_{0}^{1}\beta_{f,xy}\|\widehat{y}^{*}_{Z}(tx+(1-t)x^{\prime})-\widehat{y}^{*}_{Z^{\prime}}(tx+(1-t)x^{\prime})\|dt$
$\leq\int_{0}^{1}\beta_{f,xy}\frac{2L_{g,y}}{n\mu_{g}}dt=\frac{2L_{g,y}\beta_{f,xy}}{n\mu_{g}}.$
@@ B.2 Generalization error of the regularized exponential mechanism for bilevel SCO
Another advantage of the Regularized Exponential Mechanism is that it can have a good generalization error.
Lemma B.7.
@@
$\mathbb{E}_{x\sim\pi_{Z},Z\sim P^{n}}[\Phi(x)-\Phi^{*}]$ $\lesssim\frac{d_{x}}{k}+\mu D_{x}^{2}+\frac{L_{f,y}L_{g,y}}{\mu_{g}n}+\frac{L}{\mu}\left(\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+L_{f,x}\right)+\frac{L_{f,y}}{\mu_{g}\sqrt{k\mu}}\beta_{g,xy}(1+\kappa_{g})$
$\;\;\;+\frac{L_{f,y}CL\sqrt{n}}{\mu_{g}\mu}+\frac{L_{f,y}L_{g,y}(1+\kappa_{g})}{\mu_{g}\sqrt{n}}$
where
$\kappa_{g}:=\beta_{g,yy}/\mu_{g}$ (13)
$L:=\frac{2}{n}\left(L_{f,x}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+\frac{L_{g,y}\beta_{g,xy}}{\mu_{g}}\right)$ (14)
$C:=\beta_{g,xy}(1+\kappa_{g}).$ (15)
Proof.
We have
$\mathbb{E}_{x\sim\pi_{Z},Z\sim P^{n}}[\Phi(x)-\Phi^{*}]=\mathbb{E}_{x\sim\pi_{Z},Z\sim P^{n}}[\Phi(x)-\widehat{\Phi}_{Z}(x)+\widehat{\Phi}_{Z}(x)-\widehat{\Phi}_{Z}^{*}].$
By Lemma B.2, we have
$\mathbb{E}_{x\sim\pi_{Z},Z\sim P^{n}}[\widehat{\Phi}_{Z}(x)-\widehat{\Phi}_{Z}^{*}]\leq\frac{d_{x}}{k}+\mu D_{x}^{2}.$
Next we bound the generalization error. Note that
$\mathbb{E}_{x\sim\pi_{Z},Z\sim P^{n}}[\Phi(x)-\widehat{\Phi}_{Z}(x)]=\mathbb{E}_{x\sim\pi_{Z},Z\sim P^{n}}[F(x,y^{*}(x))-\widehat{F}_{Z}(x,\widehat{y}^{*}_{Z}(x))].$
Recall that $Z=\{z_{1},\cdots,z_{n}\}$. Suppose we replace $z_{i}\in Z$ by a fresh independent sample $z_{i}^{\prime}$ from $P$ and get $Z^{\prime}$. Then we have
$\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[F(x,y^{*}(x))-\widehat{F}_{Z}(x,\widehat{y}^{*}_{Z}(x))]$
$=$ $\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[f(x,y^{*}(x),z_{i}^{\prime})-f(x,\widehat{y}^{*}_{Z}(x),z_{i})]$
$=$ $\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x^{\prime},y^{*}(x^{\prime}),z_{i})]-\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[f(x,\widehat{y}^{*}_{Z}(x),z_{i})]$
$=$ $\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x^{\prime},y^{*}(x^{\prime}),z_{i})-f(x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(x^{\prime}),z_{i})]$
$~+\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(x^{\prime}),z_{i})]-\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x^{\prime},\widehat{y}^{*}_{Z}(x),z_{i})]$
$~+\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}),z_{i})]-\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[f(x,\widehat{y}^{*}_{Z}(x),z_{i})],$
where the first equality follows by the argument in [12, Lemma 7]. Recall that in the proof of Theorem 3.2, we show that for any $x$,
which leads to that
$\mathbb{E}_{x\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x,\widehat{y}^{*}_{Z^{\prime}}(x),z_{i})]-\mathbb{E}_{x\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x,\widehat{y}^{*}_{Z}(x),z_{i})]\leq\frac{2L_{f,y}L_{g,y}}{\mu_{g}n}.$
By Lemma B.5, we know for any $Z,z_{i}$, we have
$|f(x,\widehat{y}^{*}_{Z}(x),z_{i})-f(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}),z_{i})|=$ $|f(x,\widehat{y}^{*}_{Z}(x),z_{i})-f(x,\widehat{y}^{*}_{Z}(x^{\prime}),z_{i})+f(x,\widehat{y}^{*}_{Z}(x^{\prime}),z_{i})-f(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}),z_{i})|$
$\leq$ $\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}\|x-x^{\prime}\|+L_{f,x}\cdot\|x-x^{\prime}\|=\left(\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+L_{f,x}\right)\|x-x^{\prime}\|.$
Moreover, by Lemma B.6 and [24], we can show that $W_{2}(\pi_{Z},\pi_{Z^{\prime}})\leq\frac{L}{\mu}$ with $L=2(\frac{L_{f,x}}{n}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}n}+\frac{L_{g,y}\beta_{f,xy}}{n\mu_{g}})$. Hence we have
$\mathbb{E}_{x\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x,\widehat{y}^{*}_{Z}(x),z_{i})]-\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[f(x,\widehat{y}^{*}_{Z}(x),z_{i})]$
$=$ $\mathbb{E}_{Z,z_{i}^{\prime}}[\mathbb{E}_{x\sim\pi_{Z^{\prime}}}h_{Z,z_{i}}(x)-\mathbb{E}_{x\sim\pi_{Z}}h_{Z,z_{i}}(x)]$
$\leq$ $(\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+L_{f,y})\cdot W_{2}(\pi_{Z},\pi_{Z^{\prime}})$
$\leq$ $(\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+L_{f,y})L/\mu.$
Now, by Lipschitz continuity, it remains to bound the right-hand side of
$\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[f(x^{\prime},y^{*}(x),z_{i})-f(x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(x),z_{i})]$ $=\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[f(x,y^{*}(x),z_{i}^{\prime})-f(x,\widehat{y}^{*}_{Z}(x),z_{i}^{\prime})]$ (16)
$\leq L_{f,y}\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}\|y^{*}(x)-\widehat{y}^{*}_{Z}(x)\|.$ (17)
For any $x$, by the strong convexity of $g$, we have
$\langle\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\nabla_{y}G(x,y^{*}(x)),\widehat{y}^{*}_{Z}(x)-y^{*}(x)\rangle\geq\mu_{g}\|\widehat{y}^{*}_{Z}(x)-y^{*}(x)\|^{2}.$
By the first-order optimality, we know that $\nabla_{y}G(x,y^{*}(x))=0$, and hence we know
$\|\widehat{y}^{*}_{Z}(x)-y^{*}(x)\|\ \leq\|\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))\|/\mu_{g}.$
Similarly, we have
$\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))=0$
and
$\|\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))\|=\|\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))\|.$
Next, we bound
$\mathbb{E}_{x\sim\pi_{Z},Z}\|\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))\|.$
Let $v(Z,x):=\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))$. Using the law of total variance, we have
$\mathbb{E}[\|v(Z,x)\|^{2}]=\underbrace{\mathbb{E}_{Z}[\text{Var}_{x}(v(Z,x))]}_{\text{Term A}}+\underbrace{\text{Var}_{Z}(\mathbb{E}_{x}[v(Z,x)])}_{\text{Term B}}+\underbrace{\|\mathbb{E}_{x,Z}[v(Z,x)]\|^{2}}_{\text{Term C}}.$
Note that we have
$\|\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}))\|$
$\leq$ $~\|\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z}(x))\|+\|\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z}(x))-\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}))\|$
$\leq$ $~\beta_{g,xy}\|x-x^{\prime}\|+\beta_{g,yy}\|\widehat{y}^{*}_{Z}(x)-\widehat{y}^{*}_{Z}(x^{\prime})\|$
$\leq$ $~(\beta_{g,xy}+\frac{\beta_{g,yy}\beta_{g,xy}}{\mu_{g}})\|x-x^{\prime}\|,$ (18)
where we apply Lemma B.5 for the last step. Similarly, we also have
$\|\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))-\nabla_{y}\widehat{G}_{Z}(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}))\|\leq(\beta_{g,xy}+\frac{\beta_{g,yy}\beta_{g,xy}}{\mu_{g}})\|x-x^{\prime}\|.$
For Term A, for any dataset $Z$, $\pi_{Z}$ is $k\mu$-strongly log-concave, and function $v$ is Lipschitz in $x$ with Lipschitz constant $(\beta_{g,xy}+\frac{\beta_{g,yy}\beta_{g,xy}}{\mu_{g}})$ specified in Equation B.2. By Poincaré inequality, we have
$\mathrm{Var}_{x\sim\pi_{Z}}(v(Z,x))\leq\frac{1}{k\mu}\mathbb{E}[\|\nabla_{x}v(Z,x)\|^{2}]\leq\frac{1}{k\mu}(\beta_{g,xy}+\frac{\beta_{g,yy}\beta_{g,xy}}{\mu_{g}})^{2}.$
As for the Term B, let $v(Z)=\mathbb{E}_{x\sim\pi_{Z}}v(Z,x)$. By the Efron-Stein Inequality, we have
$\mathrm{Var}(v(Z))\leq\frac{1}{2}\sum_{j\in[n]}\mathbb{E}[\|v(Z)-v(Z_{j}^{\prime})\|^{2}]=\frac{n}{2}\mathbb{E}\|v(Z)-v(Z^{\prime})\|^{2},$
where $Z_{j}^{\prime}$ is the dataset with $z_{j}$ replaced by a fresh sample. Then we have
$\mathbb{E}\|v(Z)-v(Z^{\prime})\|^{2}=$ $~\mathbb{E}_{Z,z_{i}^{\prime}}\|\mathbb{E}_{x\sim\pi_{Z}}\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\mathbb{E}_{x\sim\pi_{Z^{\prime}}}\nabla_{y}G(x,\widehat{y}^{*}_{Z^{\prime}}(x))$
$~-(\mathbb{E}_{x\sim\pi_{Z}}\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))-\mathbb{E}_{x\sim\pi_{Z^{\prime}}}\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z^{\prime}}(x))\|^{2}.$
By Kantorovich-Rubinstein duality and Cauchy-Schwartz, we have
$\|\mathbb{E}_{x\sim\pi}h(x)-\mathbb{E}_{x^{\prime}\sim\pi^{\prime}}h(x^{\prime})\|\leq W_{1}(\pi,\pi^{\prime})\leq W_{2}(\pi,\pi^{\prime})$
for any $1$-Lipschitz function $h$ and any distributions $\pi,\pi^{\prime}\in L_{2}$. By Lemma B.5, the above fact implies
$\|\mathbb{E}_{x\sim\pi_{Z}}\widehat{y}^{*}_{Z}(x)-\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}}}\widehat{y}^{*}_{Z}(x^{\prime})\|\leq\frac{\beta_{g,xy}}{\mu_{g}}W_{2}(\pi_{Z},\pi_{Z^{\prime}})\leq\frac{\beta_{g,xy}}{\mu_{g}}L/\mu.$
A similar argument can be used to bound
$\|\mathbb{E}_{x\sim\pi_{Z}}\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}}}\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}))\|\leq\beta_{g,xy}W_{2}(\pi_{Z},\pi_{Z^{\prime}})\leq\beta_{g,xy}L/\mu,$
and likewise for $\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))$. Hence
$\mathbb{E}\|v(Z)-v(Z^{\prime})\|^{2}/100$ $\leq\mathbb{E}_{Z,z_{i}^{\prime}}\|\mathbb{E}_{x\sim\pi_{Z}}\nabla_{y}G(x,\widehat{y}^{*}_{Z}(x))-\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}}}\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}))\|^{2}$
$+\mathbb{E}_{Z,z_{i}^{\prime}}\|\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}}}\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}))-\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}}}\nabla_{y}G(x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(x^{\prime}))\|^{2}$
$+\mathbb{E}_{Z,z_{i}^{\prime}}\|\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}}}\nabla_{y}\widehat{G}_{Z^{\prime}}(x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(x^{\prime}))-\mathbb{E}_{x\sim\pi_{Z}}\nabla_{y}\widehat{G}_{Z^{\prime}}(x,\widehat{y}^{*}_{Z^{\prime}}(x))\|^{2}$
$+~\mathbb{E}_{Z,z_{i}^{\prime}}\|\mathbb{E}_{x\sim\pi_{Z}}\nabla_{y}\widehat{G}_{Z}(x,\widehat{y}^{*}_{Z}(x))-\mathbb{E}_{x\sim\pi_{Z}}\nabla_{y}\widehat{G}_{Z^{\prime}}(x,\widehat{y}^{*}_{Z^{\prime}}(x))\|^{2}$
$\leq\left(\beta_{g,xy}\frac{L}{\mu}\right)^{2}+\left(\beta_{g,yy}\left(\frac{\beta_{g,xy}}{\mu_{g}}\frac{L}{\mu}+\frac{L_{g,y}}{\mu_{g}n}\right)\right)^{2}$
$+\left(\left(\beta_{g,xy}+\frac{\beta_{g,xy}\beta_{g,yy}}{\mu_{g}}\right)\frac{L}{\mu}\right)^{2}+\left(\frac{L_{g,y}\beta_{g,yy}}{\mu_{g}n}\right)^{2}+\frac{L_{g,y}^{2}}{n^{2}}.$
Therefore,
$\mathrm{Var}(v(Z))\lesssim C^{2}n\left(\frac{L}{\mu}\right)^{2}+\frac{L_{g,y}^{2}(\kappa_{g}+1)^{2}}{n}$ (19)
for $C$ defined in the lemma statement.
We already bounded $\mathrm{Var}(v(Z))$. As $\mathbb{E}\|v(Z)\|^{2}=\mathrm{Var}(v(Z))+\|\mathbb{E}v(Z)\|^{2}$. It remains to bound Term C: $\|\mathbb{E}v(Z)\|^{2}$. For this, we have
$\mathbb{E}v(Z)=$ $\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[\nabla_{y}g(x,\widehat{y}^{*}_{Z}(x),z_{i}^{\prime})-\nabla_{y}g(x,\widehat{y}^{*}_{Z}(x),z_{i})]$
$=$ $\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[\nabla_{y}g(x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(x^{\prime}),z_{i}]-\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[\nabla_{y}g(x,\widehat{y}^{*}_{Z}(x),z_{i})]$
$=$ $\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[\nabla_{y}g(x^{\prime},\widehat{y}^{*}_{Z^{\prime}}(x^{\prime}),z_{i}]-\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[\nabla_{y}g(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}),z_{i}]$
$+\mathbb{E}_{x^{\prime}\sim\pi_{Z^{\prime}},Z,z_{i}^{\prime}}[\nabla_{y}g(x^{\prime},\widehat{y}^{*}_{Z}(x^{\prime}),z_{i}]-\mathbb{E}_{x\sim\pi_{Z},Z,z_{i}^{\prime}}[\nabla_{y}g(x,\widehat{y}^{*}_{Z}(x),z_{i})].$
Hence
$\|\mathbb{E}v(Z)\|$ $\leq\beta_{g,yy}\frac{2L_{g,y}}{\mu_{g}n}+(\beta_{g,xy}+\frac{\beta_{g,yy}\beta_{g,xy}}{\mu_{g}})W_{2}(\pi_{Z},\pi_{Z^{\prime}})$
$\leq\beta_{g,yy}\frac{2L_{g,y}}{\mu_{g}n}+(\beta_{g,xy}+\frac{\beta_{g,yy}\beta_{g,xy}}{\mu_{g}})\frac{L}{\mu}.$
Taking square roots and combining all the pieces above yields the lemma. ∎
Theorem B.8 (Precise Statement of Theorem 3.3).
@@
Grant Assumption 2.1 and parts 3 and 4 of Assumption 2.2. Assume $\widehat{\Phi}_{Z}$ and $\Phi$ are convex for all $Z$. Sampling $\widehat{x}$ from a distribution proportional to $\exp(-k(\widehat{\Phi}_{Z}(x)+\mu\|x\|^{2}/2))$ with $k=O\left(\frac{\mu n^{2}\varepsilon^{2}}{G^{2}\log(1/\delta)}\right)$ and $G=(L_{f,x}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}+\frac{L_{g,y}\beta_{f,xy}}{\mu_{g}})$ is $(\varepsilon,\delta)$-DP. Moreover,
@@
  • setting $\mu=\frac{G\sqrt{d_{x}\log(1/\delta)}}{nD_{x}\varepsilon}$, we achieve excess risk
@@
  • setting
    $\mu=\frac{1}{D_{x}}\sqrt{\frac{G}{\varepsilon n}\frac{L_{f,y}C}{\mu_{g}}+G^{2}\frac{d_{x}\log(1/\delta)}{\varepsilon^{2}n^{2}}+\frac{GCL_{f,y}}{\mu_{g}\sqrt{n}}+\frac{G}{n}\left(L_{f,x}+\frac{L_{f,y}\beta_{g,xy}}{\mu_{g}}\right)},$
    for $C:=\beta_{g,xy}(1+\kappa_{g})$ with $\kappa_{g}=\beta_{g,yy}/\mu_{g}$, the population loss has the following guarantee:
    $\mathbb{E}\Phi(\widehat{x})-\Phi^{*}$ $\lesssim D_{x}\Bigg[G\frac{\sqrt{d_{x}\log(1/\delta)}}{\varepsilon n}+\frac{1}{\sqrt{n}}\left(\sqrt{\frac{GL_{f,y}C}{\mu_{g}\varepsilon}}+\sqrt{G(L_{f,x}+L_{f,y}\beta_{g,xy}/\mu_{g})}+L_{g,y}(\kappa_{g}+1)\frac{L_{f,y}}{\mu_{g}}\right)$
    $\;\;\;+\frac{1}{n^{1/4}}\sqrt{\frac{GCL_{f,y}}{\mu_{g}}}+\frac{L_{f,y}L_{g,y}\kappa_{g}}{\mu_{g}n}\Bigg]$
@@
The population loss guarantee follows from plugging the prescribed $\mu$ and $k$ into Lemma B.7. ∎